Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time will John catch up with Martina?
Answer by checkley71(8403)

24X=48(X-2)
24X=48X-96
24X-48X=-96
-24X=-96
X=-96/-24
X=4 HOURS JOHN WILL CATCH UP WITH MARTINA.